OpenStudy (anonymous):

PHYSICS!! ok different subject but im really stuck on this answer! Which statement best compares potential and kinetic energy? A) Objects always have more potential energy than kinetic energy. B)Only kinetic energy increases when the velocity of an object increases. C)Only potential energy decreases when an object’s height increases. D)Objects always have more kinetic energy than potential energ

OpenStudy (astrophysics):

Total energy (mechanical) is \[E = T+U \implies E = \frac{ 1 }{ 2 }mv^2+mgh\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am thinking A

OpenStudy (astrophysics):

So we can see that kinetic energy is the energy of an object in motion, where potential energy is the "stored" energy, hence potential. It's not a question about which has more energy, because that doesn't really make sense, so we can eliminate A and D, because if you think of a roller coaster.. that's always a good example we have  notice that at the top it has maximum potential and 0 kinetic energy, but by the time it reaches the bottom we have maximum kinetic energy and 0 potential.